Karl and Eddie's ages added together are equal to 65. Ten years ago, Karl was twice as old as Eddie. How old are they now?
Work:
X+y= 65
(X-10)=2y-10

Answer by ikleyn(52788)

.
Karl and Eddie's ages added together are equal to 65. Ten years ago, Karl was twice as old as Eddie. How old are they now?
Work:
X+y= 65
(X-10)=2y-10
----------------------------------------------

Wrong. It should be

x + y= 65
x - 10 =2*(y - 10)